# AIDE: Antithetical, Intent-based, and Diverse Example-Based Explanations

### v1

@@ Introduction
Continuity.. We further assess the continuity metric, which refers to how well explanations capture the model behavior. Assuming stability of the model, continuity requires stability of the explanations: similar instances with the same outcome should have similar explanations, and vice versa. Sample similarity is computed using cosine similarity of embeddings, and explanation similarity is computed using Fuzzy Jaccard (Petković et al. 2021). For each sample prediction, a set is formed with the indices of training samples returned in the explanation. Fuzzy Jaccard involves solving a maximum bipartite matching problem. In spam classification, 100 random test samples are chosen. For each, the 10 most similar and dissimilar samples are identified, resulting in 2000 pairs. The same procedure is replicated with the image dataset, commencing with 50 random samples instead of 100, as this dataset is smaller in scale. The cosine similarity is plotted against Fuzzy Jaccard along with a linear regression line in red, and the Pearson correlation coefficient (PCC) for the spam datasets in Figure 2, the figures for the image dataset exhibiting the same trend can be found in the appendix. RelatIF and AIDE perform similarly. In contrast, IF and Datamodels have a lower PCC and do not exhibit a clear separation between instance pairs of low and high similarity. This is because their explanations tend to include training data outliers that appear in all explanations (globally influential), and which inflate the explanation similarity even for dissimilar pairs. Finally, TraceIn performs poorly and provides identical explanations for dissimilar points due to its extremely high susceptibility to outliers. RelatIF and AIDE are more robust because they seek to eliminate outliers, albeit in different ways (based on loss and proximity, respectively).
@@
When examining a wrong prediction for the test sample depicted in Figure 4, where the ground truth label is fish but the model predicted it as a dog, AIDE generates all four sets. After analyzing the supporters and opposers, a notable observation is the consistent presence of humans in each example. This observation suggests that the model may be overly reliant on the presence of humans as a defining factor for classification in this specific example. Furthermore, when comparing the supporters and supporters by contrast, it becomes apparent that the presence of humans serves as a key distinguishing aspect for the model to classify dogs, which is not the case when classifying fish. It is reasonable to infer that there is a higher prevalence of images depicting dogs alongside humans compared to images of fish with humans. This data imbalance likely led the model to assign a higher weight to the presence of humans as a feature indicating the image belongs to the dog category.
To confirm this, we examined test images where the model’s prediction differed from the ground truth and investigated if there was a higher presence of fish images containing humans. As expected, the images in Figure 5 were also misclassified due to this factor. Note that although the explanation of IF can also indicate the importance of humans, it does not comprehensively back up this assumption with contrastivity and by opposers who also contain humans but do not rely on it as much.
@@
When faced with an ambiguous sample as in Figure 6, where the image contains both a dog and a fish, understanding why the model chose a specific class (in this case, a dog) despite the ground truth being a fish becomes crucial. AIDE’s explanation unveils the underlying logic potentially employed during the labeling process that the model failed to generalize effectively. By examining the supporters, we observe that the model learns from both dog-related features and water-related features, which aligns with common sense. However, the opposers suggest the potential existence of a labeling rule that associates images containing both dogs and fish with the “fish” label. This rule may not have been strongly represented in the training data, leading to the model’s inefficient learning of this specific rule. Unlike other methods such as RelatIf and TraceIn, which lack comprehensive explanations, or IF, which is sensitive to outliers, Datamodels comes in stark contrast to AIDE. We observed that when confronted with mislabeled or ambiguous samples, Datamodels may explain the opposite label prediction rather than the model’s actual prediction. This happens due to a discordance between the model being explained and the intermediary models (of the same class) used to compute the importance of individual training examples; in fact, about 20% of the intermediary models predict a different class that the actual model.
Figure 6: Explanations to clarify an ambiguous prediction.
Table 3: AIDE for an ambiguous test message.
@@
Table 3 shows another ambiguous test sample for spam classification. Determining whether this message is spam or not is challenging since it does not exhibit the typical form of either a ham message or a common spam message. Instead, it takes the form of an aphorism, which falls into an ambiguous category of messages. AIDE’s supporters shed light on the presence of numerous aphorisms in the training set that are labeled as spam, indicating the existence of a labeling logic for categorizing such messages as spam. Thus, the model can correctly classify this message despite its ambiguity. The supporting samples provided by AIDE emphasize a specific logic that was likely injected during the labeling process, indicating that aphorisms were considered spam. These supporting samples contributed to the correct classification decision by reinforcing this logic. The opposing examples suggest that classifying the message as non-spam could be a plausible interpretation.
However, the model’s ability to correctly classify the message indicates that the rule regarding aphorisms being classified as spam is supported by an adequate number of training samples. This indicates that the model has learned and generalized this rule effectively.
@@
Following the recommendations of (Rong et al. 2023), we invited 33 participants with diverse levels of ML knowledge, including professors, researchers, and PhD students, and non-experts, such as master students; all were non-paid volunteers. We asked participants to classify themselves into four levels of experience with ML (Expert, Advanced, Intermediate, Beginner), which we later partition into two levels (Advanced, Intermediate). The assessment considers:
Mental Model: Q1. The explanation helped to understand the model’s prediction. To what extent do you agree?
Clarity: Q2. The explanation is clear and easy to comprehend. To what extent do you agree?
Usefulness of AIDE Quadrants: Q3, Q4, Q5, Q6. The group “S”, “SC”, “O”, “OC” enhances understanding the model’s prediction. To what extent do you agree?
Human-AI Collaboration: Q7. Did the explanation help understand how the model’s performance can be improved?
Effectiveness: Q8. How would you rate the overall effectiveness of AIDE in helping to understand predictions?
Helpfulness: Q9. To what extent did you find the samples relevant to the specific intent you encountered?
Contrastivity: Q10. Do you believe that the use of contrast in the groups of images shown enhanced your understanding of the model predictions?
@@
All questions were accompanied by a 5-point Likert scale. All positive (i.e., strongly agree, somewhat agree) answers are considered in agreement. A detailed description of the user study can be found in the appendix. The metrics collectively provide a comprehensive qualitative assessment of AIDE’s performance from the user’s perspective, taking into account various aspects of interpretability and usability. In Table 4, the percentage of participants who agreed on the high quality of specific aspects of AIDEs’ explanation for particular intents is presented. Whereas, in Table 5, the percentages of users who overall highly assessed AIDE’s effectiveness, the utility of contrast in explanation, and AIDE’s capability to tailor explanations according to user intent are depicted. A noteworthy observation is that participants with more advanced expertise tend to rate highly more frequently across various aspects of AIDE’s explanation. We also note that a positive response to the intent-based nature of explanations (Q7) can facilitate improved human-XAI collaboration, which is currently suboptimal (Schemmer et al. 2022).
Note that we do not compare AIDE with other methods to avoid participant bias, where the participant’s behavior is affected once they deduce what the preferred answers of the researcher are. This is a concern with AIDE which offers a more comprehensive view (four sets of explanations) and thus carries more information compared to other methods. Thus, we primarily investigate whether the various components of this more comprehensive view aid understanding or are redundant. Specifically, we implicitly draw conclusions on the added value of AIDE, by assessing: (1) the significance of the other three quadrants (Q4, Q5, Q6), where 63%–81% of participants agree; (2) intent nature of explanations, where 87% of participants liked; and (3) the importance of contrastivity (Q10), where 100% of the participants agree.
Table 4: Percentage (%) of people who agree with the statement on the questions Q1–7
@@
Table 5: Percentage (%) of people who agree with the statement on the questions Q8–10

### v2

@@ Introduction
Continuity. We further assess the continuity metric, which refers to how well explanations capture the model behavior. Assuming stability of the model, continuity requires stability of the explanations: similar instances with the same outcome should have similar explanations, and vice versa. Sample similarity is computed using cosine similarity of embeddings, and explanation similarity is computed using Fuzzy Jaccard (Petković et al. 2021). For each sample prediction, a set is formed with the indices of training samples returned in the explanation. Fuzzy Jaccard involves solving a maximum bipartite matching problem. In spam classification, 100 random test samples are chosen. For each, the 10 most similar and dissimilar samples are identified, resulting in 2000 pairs. The same procedure is replicated with the image dataset, commencing with 50 random samples instead of 100, as this dataset is smaller in scale. The cosine similarity is plotted against Fuzzy Jaccard along with a linear regression line in red, and the Pearson correlation coefficient (PCC) for the spam datasets in Figure 3, the figures for the image dataset exhibiting the same trend are shown in Figure 4. RelatIF and AIDE perform similarly. In contrast, IF and Datamodels have a lower PCC and do not exhibit a clear separation between instance pairs of low and high similarity. This is because their explanations tend to include training data outliers that appear in all explanations (globally influential), and which inflate the explanation similarity even for dissimilar pairs. Finally, TraceIn performs poorly and provides identical explanations for dissimilar points due to its extremely high susceptibility to outliers. RelatIF and AIDE are more robust because they eliminate outliers.
@@
When examining a wrong prediction for the test sample depicted in Figure 6, where the ground truth label is fish but the model predicted it as a dog, AIDE generates all four sets. After analyzing the S and O, a notable observation is the consistent presence of humans in each example. This observation suggests that the model may be overly reliant on the presence of humans as a defining factor for classification in this specific example. Furthermore, when comparing the S and SC, it becomes apparent that the presence of humans serves as a key distinguishing aspect for the model to classify dogs, which is not the case when classifying fish. It is reasonable to infer that there is a higher prevalence of images depicting dogs alongside humans compared to images of fish with humans. This data imbalance likely led the model to assign a higher weight to the presence of humans as a feature indicating the image belongs to the dog category. To confirm this, we examined test images where the model’s prediction differed from the ground truth and investigated if there was a higher presence of fish images containing humans. As expected, the images in Figure 7 were also misclassified due to this factor. Note that although the explanation of IF can also indicate the importance of humans, it does not comprehensively back up this assumption with contrastivity and by opposers who also contain humans but do not rely on it as much.
@@
When faced with an ambiguous sample as in Figure 2, where the image contains both a dog and a fish, understanding why the model chose a specific class (in this case, a dog) despite the ground truth being a fish becomes crucial. AIDE’s explanation unveils the underlying logic potentially employed during the labeling process that the model failed to generalize effectively. By examining the S, we observe that the model learns from both dog-related features and water-related features, which aligns with common sense. However, the O suggest the potential existence of a labeling rule that associates images containing both dogs and fish with the “fish” label. This rule may not have been strongly represented in the training data, leading to the model’s inefficient learning of this specific rule. Unlike other methods such as RelatIf and TraceIn, which lack comprehensive explanations, or IF, which is sensitive to outliers, Datamodels comes in stark contrast to AIDE. We observed that when confronted with mislabeled or ambiguous samples, Datamodels may explain the opposite label prediction rather than the model’s actual prediction. This happens due to a discordance between the model being explained and the intermediary models used to compute the importance of individual training examples; in fact, about 20% of the intermediary models predict a different class that the actual model.
Table 3: AIDE for an ambiguous text message.
@@
Table 3 shows another ambiguous test sample for spam classification. Determining whether this message is spam or not is challenging since it does not exhibit the typical form of either a ham message or a common spam message. Instead, it takes the form of an aphorism, which falls into an ambiguous category of messages. AIDE’s supporters shed light on the presence of numerous aphorisms in the training set that are labeled as spam, indicating the existence of a labeling logic for categorizing such messages as spam. Thus, the model can correctly classify this message despite its ambiguity. The supporting samples provided by AIDE emphasize a specific logic that was likely injected during the labeling process, indicating that aphorisms were considered spam. These supporting samples contributed to the correct classification decision by reinforcing this logic. The opposing examples suggest that classifying the message as non-spam could be a plausible interpretation. However, the model’s ability to correctly classify the message indicates that the rule regarding aphorisms is supported by an adequate number of training samples. This indicates that the model has learned and generalized this rule effectively.
@@
User Study
Following the recommendations of (Rong et al. 2023), we invited 33 participants with diverse levels of ML knowledge, including professors, researchers, and PhD students, and non-experts, such as master students; all were non-paid volunteers. Initial inquiries were designed to ascertain participants’ knowledge levels, facilitating subsequent segregated assessments of AIDE’s effectiveness tailored to different groups. The following questions were asked: Q: What is your level of familiarity with machine learning concepts? (Expert, Advanced, Intermediate, Beginner, Unfamiliar) Q: Are you familiar with the concept of Explainable Artificial Intelligence (XAI)? (Yes, No)
After knowing their level, we gave a detailed explanation of the dataset, task, and how the explainability technique works. A sample was shown to describe what will be presented, see Figure 8. The user study as it was presented to participants can be found here  [^9] https://forms.gle/AJKe87T9bCrqxsA47
Figure 8: A template of how explanations are presented
The assessment considers:
Mental Model: Q1. The explanation helped to understand the model’s prediction. To what extent do you agree? Clarity: Q2. The explanation is clear and easy to comprehend. To what extent do you agree? Usefulness of AIDE Quadrants: Q3, Q4, Q5, Q6. The group “S”, “SC”, “O”, “OC” enhances understanding the model’s prediction. To what extent do you agree? Human-AI Collaboration: Q7. Did the explanation help understand how the model’s performance can be improved? Effectiveness: Q8. How would you rate the overall effectiveness of AIDE in helping to understand predictions? Helpfulness: Q9. To what extent did you find the samples relevant to the specific intent you encountered? Contrastivity: Q10. Do you believe that the use of contrast in the groups of images shown enhanced your understanding of the model predictions?
The evaluation process involved separate analyses of explanations for distinct intents, aiming to gauge the alignment of AIDE with varied user needs. In the forms, we added some extra instructions to better understand what each quadrant wants to say. The study commenced with the examination of interpreting a correct prediction, illustrated in Figure 9. Subsequently, for the investigation of a wrong prediction and clarification of an ambiguous one, Figures 10 and 11 were presented, respectively. Sequentially, the questions outlined in the main paper were posed for each intent. The study concludes with general inquiries to elicit an overarching evaluation of the approach.
Figure 9: Correct prediction analysis
Figure 10: Wrong prediction analysis
Figure 11: Ambiguous prediction analysis
In Table 4 and Table 5, the results are presented through two distinct aggregation approaches. Firstly, responses from participants are combined based on their expertise level, forming two groups: “advanced”, which adds up responses from experts and advanced participants, and “intermediate”, which combines responses from beginner and intermediate participants. The second aggregation method involves the transformation of a 5-point Likert scale into a simplified two-point scale. For instance, responses such as “Agree” and “Somewhat agree” are merged into the category of “Agree”, while responses including “Disagree”, “Somewhat disagree”, and “Neutral” are consolidated into the category of “Disagree”. The subsequent discussion focuses on the results without these aggregation techniques and question by question.
Out of 33 participants, the distribution across expertise levels was 3 experts, 13 advanced, 12 intermediate, and 5 beginners. The possible answers were given in a 5-point Likert scale, denoted as follows: OO – Strongly agree; O – Somewhat agree; N – Neither agree nor disagree; X – Somewhat disagree; XX – Disagree.
@@
In Table 4, the percentage of participants who agreed on the high quality of specific aspects of AIDEs’ explanation for particular intents is presented. Whereas, in Table 5, the percentages of users who overall highly assessed AIDE’s effectiveness, the utility of contrast in explanation, and AIDE’s capability to tailor explanations according to user intent are depicted. Notably, participants with more advanced expertise tend to rate highly more frequently across various aspects of AIDE’s explanation. We also note that a positive response to the intent-based nature of explanations (Q7) can facilitate improved human-XAI collaboration, which is currently suboptimal (Schemmer et al. 2022).
Table 4: Percentage (%) of people who agree with Q1–7.
@@
Table 5: Percentage (%) of people who agree with Q8–10.
@@
Note that we do not compare AIDE with other methods to avoid participant bias, where the participant’s behavior is affected once they deduce what the preferred answers of the researcher are. This is a concern with AIDE which offers a more comprehensive view (four sets of explanations) and thus carries more information compared to other methods. Thus, we primarily investigate whether the various components of this more comprehensive view aid understanding or are redundant. Specifically, we implicitly draw conclusions on the added value of AIDE, by assessing: (1) the significance of the other three quadrants (Q4, Q5, Q6), where 63%–81% of participants agree; (2) intent nature of explanations, where 87% of participants liked; and (3) the importance of contrastivity (Q10), where 100% agree.
Other findings from the user study can be found in the following tables.
Intent Level OO O N X XX
Expert 1 2
Advanced 8 5
Intermediate 8 2 1 1
Int. correct Beginner 2 3
Expert 1 2
Advanced 4 7 1 1
Intermediate 4 3 2 3
Inv. wrong Beginner 1 4
Expert 1 1 1
Advanced 7 2 3 1
Intermediate 4 5 1 2
Cl. ambiguous Beginner 2 2 1
Table 6: Mental Model: Q1 – The explanation provided helped to understand the model’s prediction. To what extent do you agree?
Intent Level OO O N X XX
Expert 2 1
Advanced 9 3 1
Intermediate 7 3 2
Int. correct Beginner 2 3
Expert 2 1
Advanced 4 7 1 1
Intermediate 4 3 2 3
Inv. wrong Beginner 2 3
Expert 2 1
Advanced 7 4 2
Intermediate 4 5 1 2
Cl. ambiguous Beginner 2 2 1
Table 7: Clarity: Q2 – The explanation is clear and easy to comprehend. To what extent do you agree?
Question Level OO O N X XX
Expert 4 3 1 1
Advanced 23 14 1 1
Intermediate 22 9 2 3
Q3 Beginner 8 6 1 1
Expert 2 2 2
Advanced 13 8 3 1 1
Intermediate 12 4 2 5 1
Q4 Beginner 5 3 1 1
Expert 2 1 1 2
Advanced 7 8 7 4
Intermediate 10 8 2 4
Q5 Beginner 3 6 1
Expert 2 1
Advanced 3 5 4 1
Intermediate 4 5 1 2
Q6 Beginner 2 6 1 3
Table 8: Usefulness of AIDE Quadrants: Q3/Q4/Q5/Q6 – The group “Support”/“Support by Contrast”/“Oppose”/“Oppose by Contrast” enhances the understanding of the model’s prediction. To what extent do you agree?
Intent Level Yes No
Expert 1 1
Advanced 12 1
Intermediate 11 1
Inv. wrong Beginner 4 1
Expert 2 1
Advanced 9 4
Intermediate 9 3
Cl. ambiguous Beginner 3 2
Table 9: Human-AI Collaboration: Q7 – Did the explanation help understand how the model’s performance can be improved?
Question Level OO O N X XX
Expert 1 2
Advanced 4 7 2
Intermediate 2 8 1 1
Q8 Beginner 2 3
Expert 3
Advanced 6 7
Intermediate 4 6 2
Q9 Beginner 1 2 2
Expert 1 2
Advanced 9 4
Intermediate 3 8 1
Q10 Beginner 4 1
Table 10: Effectiveness: Q8 – How would you rate the overall effectiveness of AIDE in helping to understand model predictions?; Helpfulness: Q9 – To what extent did you find the provided samples relevant to the specific intent you encountered?; Contrastivity: Q10 – Do you believe that the use of contrast in the groups of images shown enhanced your understanding of the model predictions?
@@ Conclusion
Limitations
One limitation is that we exclusively tested the approach on binary classification. Extending our work beyond binary classification mostly concerns the ”opposite label-same label” dimension. There are two options: (a) one-vs-rest, and group all opposite labels together similar to the ”by contrast” quadrants, or (b) one-vs-one, and break down opposite labels. This impacts presentation where the latter approach requires users to build a bit more complex mental model.
Another limitation is our user study’s relatively small sample size, comprising only 33 participants with diverse backgrounds. This limited sample size precluded the possibility of conducting a large-scale and extensive survey, where we could include examples from existing methods for a control group not exposed to the AIDE explanations, thus mitigating the bias discussed above. By integrating results from both groups, we could more robustly demonstrate the superiority of our methods compared to the baselines.
Acknowledgments
This work was supported by the European Union under Horizon 2020 grant agreement No. 955895.